One mole of a diatomic ideal gas is cooled from 200oC to 50oC. Determine the change in the translational kinetic energy and the change in the internal energy of the gas
1
Expert's answer
2019-06-04T12:07:28-0400

The change in the translational kinetic energy:

"\Delta K=\frac{3}{2}nR(T_2-T_1)"

"\Delta K=\frac{3}{2}(1)(8.31)(50-200)=-1870\ J"

The change in the internal energy of the gas:


"\Delta U=\frac{5}{2}nR(T_2-T_1)"

"\Delta U=\frac{5}{2}(1)(8.31)(50-200)=-3120\ J"